# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

### v2

@@ 3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
We propose an adaptive-rank algorithm for the GSE (13). We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format, i.e.:
[…]
where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times. Projection methods involve constructing bases $U_{1}$ and $V_{1}$ and then projecting the original equation onto them to determine the coefficients of $S_{1}$. A key aspect of the proposed adaptive-rank approach is that we never form the full solutions explicitly. Instead, we seek bases in a dimension-by-dimension manner and then use them to perform a Galerkin projection. Key questions for solving the GSE that we address in this study include:
  • How do we choose suitable subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
  • How can we solve the reduced equation for $S_{1}$ efficiently?
  • How can we check the residual of the GSE without forming the matrices explicitly?
Our main innovation leverages an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (12). This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning for solving the projected reduced system for $S_{1}$, in the spirit of [31]. In the following, we cover the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4. Finally, Section 3.5 extends the algorithm to the case of multi-rank coefficients.
@@ 3.1 Averaged-Coefficient Approximate Operator $\tilde{\mathscr{L}}$
Our approach incorporates ideas from [31] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework. The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
@@
~${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (16)
@@
$A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\quad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\quad P_{1}=A_{1}+A_{3},$ (18)
$A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\quad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\quad P_{2}=A_{2}+A_{4}.$ (19)
The solution to (17) is an approximation to that of (13), but it can be efficiently computed using low-rank approximations [14, 39, 37, 36]. It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the GSE (13). This was done in [31] with $\mathcal{O}(N^{3})$ computational complexity. In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the GSE with GMRES. This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$.
@@ 3.2 Basis Construction
We begin by defining the xKrylov subspaces used for our basis construction at the $m$-th outer iteration as:
$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big{\{}U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big{\}}.$ (20)
where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our GSE, and $l$ is the number of operators used for each basis augmentation step.
The xKrylov basis construction involves several operators from the original GSE (13) and its approximation (17). Following [14], we construct an xKrylov subspace in the $x$-direction for the approximated SE (17) using $P_{1}$ and its inverse. We also include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$. Additionally, we consider the action of the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$. We have found that including all these matrices is effective in capturing the range of the solution. Using these operators, we construct a generalized Krylov subspace in the $x$-direction as follows:
$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\quad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
Depending on the structure of the GSE (2), $l$ may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies. To limit subspace inflation, we employ an SVD-based truncation strategy (Algorithm 1), such that only basis vectors corresponding to significant singular values are retained.
The construction of the dimension-wise xKrylov subspaces (20), following [39], is as follows. At each iteration $m$, Algorithm 2 augments the previously constructed basis $\mathcal{V}_{(m-1)}$ by applying powers of operators $M_{1}$ through $M_{l}$ to their corresponding basis blocks $V^{(1)},\dots,V^{(l)}$. Each block matrix $V^{(i)}$ may have a rank of at most $r_{0}$. We apply Algorithm 2 to orthogonalize the subspace and obtain a basis $\mathcal{U}_{(m)}\in\mathbb{R}^{N_{1}\times r^{(m)}_{x}}$ and $\mathcal{V}_{(m)}\in\mathbb{R}^{N_{2}\times r^{(m)}_{y}}$ in the $x$ and $y$ directions, respectively. For notational simplicity, we omit the calligraphic notation and subscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration. The truncation step in line 2 of Algorithm 2 removes any linear dependencies within each block during the power iterations; line 2 applies a global truncation across all newly generated blocks, eliminating redundancies among them. As a result, the basis of the subspace is kept optimal in size.
@@
Output: Truncated bases ${Q}$.
@@ 1 Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
@@ 7 $\mathcal{V}_{(m)}\leftarrow[\mathcal{V}_{(m-1)},V^{\prime}]$;
We comment next on the performance of the xKrylov iteration. In principle, the number of xKrylov iterations may depend on the mesh size, $N$. For instance, [24, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$, with $\kappa\sim N^{2}$ the condition number. While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice. In particular, for moderately tight tolerances, we find that the required Krylov subspace dimensions remain independent of grid resolution for a wide range of grid sizes $N$.
@@ 3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-Preconditioned GMRES
${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\qquad\text{with}\qquad\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (21)
@@
$\tilde{T}_{1}$ $=U_{1}^{\top}T_{1}U_{1},\quad\tilde{T}_{3}=U_{1}^{\top}T_{3}U_{1},\quad\tilde{\Phi}_{1}=U_{1}^{\top}\Phi_{1}U_{1},\quad\tilde{\Sigma}_{1}=U_{1}^{\top}\Sigma_{1}U_{1},$
$\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\quad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\quad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\quad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (22) with a direct solver scales as $\mathcal{O}(r^{6})$ in 2D and $\mathcal{O}(r^{9})$ in 3D, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values. Instead, we solve this equation iteratively using preconditioned GMRES. Instead of adapting GMRES to the matrix equation, we wrap GMRES around the Kronecker form, but exploiting the matrix-equation form (21) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below, with overall complexity $\mathcal{O}(r^{3})$ in 2D (and $\mathcal{O}(r^{4})$ in 3D). For the GMRES operator-vector product, given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is performed as follows:
@@
    Compute $\hat{\mathbf{Z}}_{k}\leftarrow\mathbf{S}_{k}-\Delta t\left(\tilde{T}_{1}\mathbf{S}_{k}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}\mathbf{S}_{k}\tilde{T}_{2}^{\top}+\tilde{T}_{3}\mathbf{S}_{k}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}\mathbf{S}_{k}\tilde{T}_{4}^{\top}\right)$, with complexity $\mathcal{O}(r^{3})$.
@@
For efficiency, GMRES requires preconditioning so that the eigenvalues of the preconditioned matrix are sufficiently clustered to deliver $r$-independent convergence rates  [7]. Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (17). The ACS preconditioner is expected to deliver optimal rank-independent, superlinear convergence (both observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations, [^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one. which are known to be effective preconditioners for Krylov methods [1, 2].
We consider left preconditioning as:
@@
    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, for example with the Bartels-Stewart algorithm [15].
@@
3.4 Truncation, Efficient Residual Computation

### v3

@@ 3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
We describe next the projection-based framework for solving the GSE (12) [5, 20, 45, 36]. We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format, i.e.:
$\mathbf{F}_{0}=U_{0}S_{0}V_{0}^{\top}\xrightarrow{\text{evolve in time}}\mathbf{F}_{1}=U_{1}S_{1}V_{1}^{\top},$ (14)
where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times. Projection methods involve constructing bases $U_{1}$ and $V_{1}$ and then projecting the original equation onto them to determine the coefficients of $S_{1}$. Efficiently solving the GSE requires choosing suitable solution subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$, and solving the reduced equation for $S_{1}$ efficiently. To accomplish both, in the spirit of the work in [35] we leverage an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator $\mathscr{L}$} in (11). In the following, we describe the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our particular implementation of the ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4. Finally, Section 3.5 extends the algorithm to the case of multi-rank coefficients.
@@ 3.1 Averaged-Coefficient Approximate Operator $\tilde{\mathscr{L}}$
We incorporate ideas from [35] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework. The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) contribute to constructing the solution bases $U_{1}$ and $V_{1}$ from xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
@@
${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (15)
@@
$A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\qquad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\qquad P_{1}=A_{1}+A_{3},$ (17)
$A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\qquad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\qquad P_{2}=A_{2}+A_{4}.$ (18)
Equation (16) is an SE approximation to the GSE (12), which can be efficiently solved using low-rank approximations [14, 42, 40, 39]. It can be exploited as a preconditioner for a GMRES iterative solution of the GSE (12). This was done in [35] with $\mathcal{O}(N^{3})$ computational complexity. In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the GSE with GMRES. This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity.
@@ 3.2 Basis Construction
The construction of the the xKrylov subspaces per dimension at the $m$-th outer iteration is performed as:
$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big\{U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big\},$ (19)
where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in the GSE, and $l$ is the number of operators used for each basis augmentation step. For the $M_{l}$ operators, we include $P_{1}$ and its inverse from the SE approximation to the GSE (16), the individual operators $A_{1}$ and $A_{3}$ and their inverses, as they may contain information not captured by the inverse of $P_{1}$, and the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$. Using these operators, the generalized Krylov subspace (for instance in the $x$-direction) considered is as follows:
$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\qquad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
We have found that including all these matrices is important to capture the range of the solution.
Depending on the structure of the GSE (2), $l$ in (19) may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies. To limit subspace inflation, we employ an SVD-based truncation strategy (Algorithm 1), such that only basis vectors corresponding to significant singular values are retained. With this in mind, the construction of the dimension-wise xKrylov subspaces (19), following [42], is as follows. At each iteration $m$, Algorithm 2 augments the previously constructed basis $\mathcal{V}_{(m-1)}$ by applying powers of operators $M_{1}$ through $M_{l}$ to their corresponding basis blocks $V^{(1)},\dots,V^{(l)}$. Each block matrix $V^{(i)}$ may have a rank of at most $r_{0}$. We apply Algorithm 2 to orthogonalize the subspace and obtain a basis $\mathcal{U}_{(m)}\in\mathbb{R}^{N_{1}\times r^{(m)}_{x}}$ and $\mathcal{V}_{(m)}\in\mathbb{R}^{N_{2}\times r^{(m)}_{y}}$ in the $x$ and $y$ directions, respectively. For notational simplicity, we omit the calligraphic notation and subscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration. The truncation step in line 2 of Algorithm 2 removes any linear dependencies within each block during the power iterations; line 2 applies a global truncation across all newly generated blocks, eliminating redundancies among them. As a result, the subspace is kept optimal in size.
@@
Output: Truncated bases Q{Q}.
@@ 1 Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ ;
Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$ ;
// Optional: Set Maximum Allowable Rank
@@ 7 $\mathcal{V}_{(m)}\leftarrow[\mathcal{V}_{(m-1)},V^{\prime}]$;
We comment next on the expected performance of the xKrylov iteration. In principle, the number of xKrylov iterations may depend on the mesh size, $N$. For instance, [25, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$ in 2D, and $(\kappa^{2/3})\sim\mathcal{O}(N^{4/3})$ in 3D, with $\kappa\sim N^{2}$ the condition number. While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice, with Krylov subspace dimensions remaining largely independent of grid resolution for moderate convergence tolerances and mesh refinements.
@@ 3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-Preconditioned GMRES
${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\hskip 18.49988pt\text{with}\hskip 18.49988pt\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (20)
@@
$\tilde{T}_{1}$ $=U_{1}^{\top}T_{1}U_{1},\qquad\tilde{T}_{3}=U_{1}^{\top}T_{3}U_{1},\qquad\tilde{\Phi}_{1}=U_{1}^{\top}\Phi_{1}U_{1},\qquad\tilde{\Sigma}_{1}=U_{1}^{\top}\Sigma_{1}U_{1},$
$\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\qquad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\qquad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\qquad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (21) with a direct solver scales as $\mathcal{O}(r^{3d})$, with $r$ the maximum rank of all dimensions and $d$ the dimensionality, which may be too expensive even for moderate $r$-values. Instead, we solve this equation iteratively using preconditioned GMRES. We wrap standard GMRES around the Kronecker form of the GSE, but exploiting the matrix-equation form (20) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below. The resulting overall complexity is $\mathcal{O}(r^{d+1})$. For the GMRES operator-vector product, given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is performed as follows:
@@
    Compute $\hat{\mathbf{Z}}_{k}\leftarrow\mathbf{S}_{k}-\Delta t\left(\tilde{T}_{1}\mathbf{S}_{k}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}\mathbf{S}_{k}\tilde{T}_{2}^{\top}+\tilde{T}_{3}\mathbf{S}_{k}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}\mathbf{S}_{k}\tilde{T}_{4}^{\top}\right)$, with complexity $\mathcal{O}(r^{d+1})$.
@@
For efficiency, GMRES requires preconditioning so that the eigenvalues of the preconditioned matrix are sufficiently clustered to deliver $r$-independent convergence rates  [7]. Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (16), with $\tilde{P}_{1}=U_{1}^{\top}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{\top}P_{2}V_{1}$. The ACS preconditioner delivers optimal, rank-independent, superlinear convergence (all observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations. [^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one, which are known to be effective preconditioners for Krylov methods [1, 2]. In particular, we consider left preconditioning as:
@@
    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, e.g. with the Bartels-Stewart algorithm [15].
@@
3.4 Truncation and Residual Computation